# Fractional Sobolev paths on Wasserstein spaces and their energy-minimizing particle representations

### v3

@@ Abstract.
We study a generalization of Kantorovich’s optimal transportation problem. Given a prescribed family of time-dependent probability measures $(\mu_{t})$, we aim to find, among all path-continuous stochastic processes whose one-dimensional time marginals coincide with $(\mu_{t})$ (if there is any), a process that minimizes a given energy. After discussing a sufficient condition for the energy to ensure the existence of a minimizer, we investigate fractional Sobolev energies. Given a deterministic path $(\mu_{t})$ on a $p$-Wasserstein space with fractional Sobolev regularity $W^{\alpha,p}$, where $11/p<\alpha<$ 1, we provide conditions under which we prove the existence of a process that minimizes the energy and construct a process that realizes the regularity of $(\mu_{t})$. While continuous paths of low regularity on Wasserstein spaces naturally appear in stochastic analysis, they can also arise deterministically as solutions to the continuity equation. This paper is devoted to the deterministic setting to gain some understanding of the required conditions.
@@ 1.1. Problem formulation and main results
Let $(\mathcal{X},d)$ be a complete separable metric space and $\mathcal{B}(\mathcal{X})$ be the $\sigma$-algebra of Borel sets of $\mathcal{X}$. Let $P(\mathcal{X})$ denote the set of Borel probability measures on $\mathcal{X}$ and $P_{p}(\mathcal{X})\subset P(\mathcal{X})$ denote the subset of measures with finite $p$-th moment for $p\geq 1$. Consider a prescribed family of probability measures $(\mu_{t})\coloneqq(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ indexed by $t$ in a time interval $I\coloneqq[0,T]\subset\mathbb{R}$. In many applications, it is of interest to have a stochastic process $(X_{t})$ on a suitable path space $\Gamma_{T}\subset\mathcal{X}^{[0,T]}$ whose one-dimensional time marginals coincide with $(\mu_{t})$. We know that this requirement does not uniquely determine the process in general. In this paper, we are interested in processes whose path laws, denoted by $\pi\in P(\Gamma_{T})$ and referred to as lifts, carry further intrinsic information about $(\mu_{t})$, as a result of a minimization problem analogous to Kantorovich’s optimal transport problem (1942) [12]. Typical choices for $\Gamma_{T}$ are the space of continuous paths and the space of càdlàg paths. Throughout this paper, we consider $\Gamma_{T}\coloneqq C([0,T];\mathcal{X})$ endowed with the $\sigma$-algebra $\mathcal{C}$ generated by the corresponding evaluation maps $\{e_{t}:t\in I\}$. We also equip $\Gamma_{T}$ with the supremum distance and recall that the Borel $\sigma$-algebra generated by the corresponding open sets coincides with $\mathcal{C}$. We formulate the problem mentioned above as a variational problem on the space of path measures on $\Gamma_{T}$:
@@
Let $\Psi:\Gamma_{T}\to[0,+\infty]$ be a measurable functional. Given $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ with $I\coloneqq[0,T]\subset\mathbb{R}$, consider the variational problem
@@
This extends Kantorovich’s classical problem, with two fixed marginals $\mu,\nu\in P(\mathcal{X})$, to a time-dependent setting with infinitely many fixed marginals $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$. In contrast to the set of couplings $\mathrm{Cpl}(\mu,\nu)\subset P(\mathcal{X}^{2})$, which is always non-empty, the set of liftings $\mathrm{Lift}(\mu_{t})$ $\subset P(\Gamma_{T})\subset P(\mathcal{X}^{I})$ can be empty. In this case, the infimum (1.1) is +$\infty$+ ∞ by the usual convention.
One can always obtain a process from a lift $\pi$ by looking at its associated canonical process. From a physical perspective, this provides a particle representation, and the objective function in 1 can be interpreted as the total cost of the particles’ motion. We thus call $\int\Psi{\mathrm{\,d}}\pi$ the energy of $\pi$ with respect to the energy functional $\Psi$. The probabilistic formulation of the problem above is given in Remark 1.12.
When $\mu,\nu\in P_{p}(\mathcal{X})$ for some $p\geq 1$ and the cost function in Kantorovich’s problem is taken as $d(x,y)^{p}$, we know an optimal coupling, denoted by $\Upsilon\in\mathrm{OptCpl}(\mu,\nu)$, exists and defines the $p$-(Kantorovitch–Rubinstein–)Wasserstein distance: $W^{p}_{p}(\mu,\nu)=\int_{\mathcal{X}^{2}}d(x,y)^{p}{\mathrm{\,d}}\Upsilon(x,y)$.
Now let $(\mu_{t})_{t\in I}\subset P_{p}(\mathcal{X})$ be a path on the Wasserstein space $(P_{p}(\mathcal{X}),W_{p})$ with a certain path regularity $|\mu|<+\infty$. Here, ||[0,+]$\cdot$|$\in$[0,$\infty$]| ⋅ | ∈ [ 0 , + ∞ ] represents a functional that captures a certain regularity of curves in metric spaces (e.g. absolute continuity or Hölder regularity). In view of the two-marginal case, one can similarly ask whether there exists a lift $\pi$ that realizes this regularity in the sense that
@@
Similarly, one expects that any alternative lift produces higher energy. This is indeed the case for the regularities previously studied and those in this paper. This observation has motivated introducing 1, where the general functional $\Psi$ is chosen here as $|\cdot|^{p}$. We call a lift minimizing or optimal if it achieves the minimum possible energy. We further call it realizing if it attains equality (1.3) (we will observe that in some cases, a minimizing lift exists but a realizing one doesn’t). Lifts that realize the regularity of Wasserstein curves have been previously constructed under different assumptions on regularity, some of which are highlighted here:
@@
    $p$-absolutely continuous curves in $p$-Wasserstein spaces with $p>1$ on separable Hilbert spaces: Ambrosio–Gigli–Savaré [4] characterize these curves via the continuity equation and obtained a vector field $v_{t}$ that realizes the metric speed i.e. $|\dot{\mu}_{t}|^{p}=\int_{\mathcal{X}}|v_{t}|^{p}{\mathrm{\,d}}\mu_{t}$ for a.e. $t\in I$.
@@
Our study of low-regularity paths on Wasserstein spaces is motivated by measure-valued solutions to conservative stochastic PDEs, such as stochastic Fokker–Planck–Kolmogorov equations. Low-regularity paths can also arise deterministically as solutions to the continuity equation, though less trivially. We focus on the deterministic setting in this paper and study the random setting in a subsequent paper.
Since, in low regularity, it is less straightforward which norm can be minimized and potentially realized, our first step is to identify a sufficient condition on the energy to guarantee the existence of a minimizer. Below is Proposition 3.1—a simple observation using the direct method in the calculus of variations.
@@
where $\bar{x}\in\mathcal{X}$ is an arbitrary point and $|\cdot|:C(I;\mathcal{X})\to[0,+\infty]$ is a (semi-)norm. When $(\mathcal{X},d)$ has enough structure so that closed bounded sets are compact, Arzelà-Ascoli theorem can verify which norms $|\gamma|$ make the sublevels of $\Psi$ relatively compact. This is summarized for some commonly used norms in low-regularity settings in Table 1 at the end of this section. We note that the modulus of continuity $w_{\delta}(\gamma)$, $p$-variation $|\gamma|_{p\textrm{-}\mathrm{var}}$, and its infinitesimal characterization $|\gamma|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}$ fail these conditions. In particular, quadratic variation $[\gamma]$ in the sense of stochastic analysis also fails; see Remark 1.13. By contrast, Hölder regularity $|\gamma|_{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$, fractional Sobolev regularity $|\gamma|_{W^{\alpha,p}}$, certain Besov regularity $|\gamma|_{b^{\alpha,p}}$, and Sobolev regularity $|\gamma|_{W^{1,p}}$ satisfy the required conditions. Definitions and properties of these norms are given in Sections 2.1, 2.2, 2.3, 2.4 and 2.5. The function space $W^{1,p}(I;\mathcal{X})$ used in this paper coincides with the set of $p$-absolutely continuous curves $AC^{p}(I;\mathcal{X})$; thus, the last item in the table is already addressed by Lisini’s results [14].
Following this observation, we consider paths in the fractional Sobolev space $W^{\alpha,p}$ or certain Besov space $b^{\alpha,p}$. These spaces also contain Hölder spaces for specific exponents. Given a metric space $(\mathcal{X},d)$, an integrability parameter $1\leq p<\infty$, and a regularity parameter $10<\alpha<$ 1, these function spaces are defined as:
@@
Let $(\mathcal{X},d)$ be a complete separable metric space in which closed bounded sets are compact, and $I\coloneqq[0,T]\subset\mathbb{R}$. Let $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ be such that $\mu_{0}\in P_{p}(\mathcal{X})$ and it has a lift with finite $W^{\alpha,p}$-energy with $1<p<\infty$ and $\frac{1}{p}<\alpha<1$. Then, there exists a minimizer $\pi\in P(C(I;\mathcal{X}))$ to 1 for the energy $\Psi(\gamma)=|\gamma|^{p}_{W^{\alpha,p}}$. In particular,
@@
    Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ can form a compatible collection under suitable conditions on their covariance matrices. See [19, page 49].
@@
    For more nontrivial compatible examples, see [7, Proposition 4.1] and [19, Section 2.3.2].
@@
Figure 1. An illustration of two different constructions, showing which two-dimensional marginals of $\Upsilon_{n}$ are optimal. Here, $n=2$ and the time interval $[0,1]$ is divided into $42^{2}=$ 4 equal pieces.
@@
[…]
@@
Another natural question is whether it is possible to remove the assumptions of the existence of a lift with finite energy and compatibility from Theorem 1.3 and Theorem 1.6, respectively, while still obtaining a minimizing lift. We demonstrate that the answer to this question can be negative. In other words, if compatibility is dropped, the infimum (1.1) for the energy $\Psi(\gamma)=|\gamma|^{p}_{W^{\alpha,p}}$ can be +$\infty$+ ∞. More specifically, in Example 4.5, we show:
@@
There exists a non-compatible curve $(\mu_{t})\in W^{\alpha,p}(I;P_{p}(\mathcal{X}))$ for some $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, whose only existing lift $\pi$ on continuous paths has infinite $W^{\alpha,p}$-energy.
@@
As a final result, we give a dynamic formulation of the Wasserstein distance using Besov energy. We first show in Lemmas 3.11 and 3.12 that on a metric space $(\mathcal{X},d)$ and for any $1<p<\infty$ and $10<\alpha<$ 1, the following are equivalent:
@@
Notice that $\left(1-2^{-(p-\alpha p)}\right)\in(0,1)$ on the parameter range above. This characterization allows us to give the result below, which is a generalization of the metric Benamou–Brenier formula [5], [4, Eq. (8.0.3)], [14, Corollary 4], to the fractional setting. Below is Corollary 3.13 and $Geo([0,1];\mathcal{X})$ denotes the set of constant-speed geodesics.
@@
Let $(\mathcal{X},d)$ be a complete, separable, and geodesic metric space. Let $1<p<\infty$ and $10<\alpha<$ 1. Then for every $\mu,\nu\in P_{p}(\mathcal{X})$, we have
@@
In other words, the task is to find a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a stochastic process $(X_{t})$ with marginals $(\mu_{t})$ such that the expectation (1.18) is minimized. As before, if $\mathrm{Process}(\mu_{t})=\varnothing$, the infimum (1.18) is set +$\infty$+ ∞. As in the classical optimal transport, only the distribution of these processes enters the cost function. Thus, we will only focus on the formulation of 1.
@@
Given $\gamma\in C([0,T];\mathcal{X})$, the stochastic calculus quadratic variation $[\gamma]_{T}$ over $[0,T]$ is usually defined by taking a limit over shrinking partitions of $[0,T]$. By the example in Remark 2.14, one sees that the map $\gamma\mapsto[\gamma]_{T}$ from $C([0,T];\mathcal{X})\to[0,+\infty]$ is not lower semi-continuous. Furthermore, in Example 4.3, by taking $\upgamma=\tfrac{1}{2}$ and $p>2$, one sees that there exists a determinist curve $(\mu_{t})\subset P_{p}(\mathcal{X})$ with non-zero quadratic variation with respect to $W_{p}$, whose only path-continuous process $(X_{t})$, equivalently its lift $\pi$, has absolutely continuous sample paths. In particular,
$[\mu]_{T}^{2}>\mathbb{E}\Big{[}[X]_{T}^{2}\Big{]}=\int_{\Gamma_{T}}[\gamma]_{T}^{2}{\mathrm{\,d}}\pi=0,$ (1.20)
which shows that $[\mu]_{T}$ cannot even be bounded by the quadratic variation energy of its lift.

### v4

@@ Abstract.
We study a generalization of the Monge–Kantorovich optimal transport problem. Given a prescribed family of time-dependent probability measures $(\mu_{t})$, we aim to find, among all path-continuous stochastic processes whose one-dimensional time marginals coincide with $(\mu_{t})$ (if there is any), a process that minimizes a given energy. After discussing a sufficient condition for the energy to ensure the existence of a minimizer, we investigate fractional Sobolev energies. Given a deterministic path $(\mu_{t})$ on a $p$-Wasserstein space with fractional Sobolev regularity $W^{\alpha,p}$, where $1/p<\alpha<1$, we provide conditions under which we prove the existence of a process that minimizes the energy and construct a process that realizes the regularity of $(\mu_{t})$. While continuous paths of low regularity on Wasserstein spaces naturally appear in stochastic analysis, they can also arise deterministically as solutions to the continuity equation. This paper is devoted to the deterministic setting to gain some understanding of the required conditions. The subsequent companion paper (arXiv:2503.10859) focuses on the stochastic setting and applications to SPDEs.
@@ 1.1. Problem formulation and main results
Let $(\mathcal{X},d)$ be a complete separable metric space and $\mathcal{B}(\mathcal{X})$ be the $\sigma$-algebra of Borel sets of $\mathcal{X}$. Let $P(\mathcal{X})$ denote the set of Borel probability measures on $\mathcal{X}$ and $P_{p}(\mathcal{X})\subset P(\mathcal{X})$ denote the subset of measures with finite $p$-th moment for $p\geq 1$. Consider a prescribed family of probability measures $(\mu_{t})\coloneqq(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ indexed by $t$ in a time interval $I\coloneqq[0,T]\subset\mathbb{R}$. In many applications, it is of interest to have a stochastic process $(X_{t})$ on a suitable path space $\Gamma_{T}\subset\mathcal{X}^{[0,T]}$ whose one-dimensional time marginals coincide with $(\mu_{t})$. We know that this requirement does not uniquely determine the process in general. In this paper, we are interested in processes whose path laws, denoted by $\pi\in P(\Gamma_{T})$ and referred to as lifts, carry further intrinsic information about $(\mu_{t})$, as a result of a minimization problem analogous to the transport problem by Kantorovich in 1942 [15], following Monge [22]. Typical choices for $\Gamma_{T}$ are the space of continuous paths and the space of càdlàg paths. Throughout this paper, we consider $\Gamma_{T}\coloneqq C([0,T];\mathcal{X})$ endowed with the $\sigma$-algebra $\mathcal{C}$ generated by the corresponding evaluation maps. We also equip $\Gamma_{T}$ with the supremum distance and recall that the Borel $\sigma$-algebra generated by the corresponding open sets coincides with $\mathcal{C}$. We formulate the problem mentioned above as a variational problem on the space of path measures on $\Gamma_{T}$:
@@
Given $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ with $I\coloneqq[0,T]\subset\mathbb{R}$ and a measurable functional $\Psi:\Gamma_{T}\to[0,+\infty]$, consider the variational problem
@@
This extends Monge–Kantorovich's classical problem, with two fixed marginals $\mu,\nu\in P(\mathcal{X})$, to a time-dependent setting with infinitely many fixed marginals $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$. In contrast to the set of couplings $\mathrm{Cpl}(\mu,\nu)\subset P(\mathcal{X}^{2})$, which is always non-empty, the set of liftings $\mathrm{Lift}(\mu_{t})$ $\subset P(\Gamma_{T})\subset P(\mathcal{X}^{I})$ can be empty. In this case, the infimum (1.1) is $\infty$ by the usual convention.
One can always obtain a process from a lift $\pi$ by looking at its associated canonical process. From a physical perspective, this provides a particle representation, and the objective function in 1 can be interpreted as the total cost of the particles' motion. We thus call $\int\Psi{\mathrm{\,d}}\pi$ the energy of $\pi$ with respect to the energy functional $\Psi$. The probabilistic formulation of the problem above is given in Remark 1.12.
When $\mu,\nu\in P_{p}(\mathcal{X})$ for some $p\geq 1$ and the cost function in Monge–Kantorovich’s problem is taken as $d(x,y)^{p}$, we know an optimal coupling, denoted by $\Upsilon\in\mathrm{OptCpl}(\mu,\nu)$, exists and defines the pp-(Kantorovitch–Rubinstein–)Wasserstein distance: $W^{p}_{p}(\mu,\nu)=\int_{\mathcal{X}^{2}}d(x,y)^{p}{\mathrm{\,d}}\Upsilon(x,y)$.
Now let $(\mu_{t})_{t\in I}\subset P_{p}(\mathcal{X})$ be a path on the Wasserstein space $(P_{p}(\mathcal{X}),W_{p})$ with a certain path regularity $|\mu|<+\infty$. Here, $|\cdot|\in[0,+\infty]$ represents a functional that captures a certain regularity of curves in metric spaces (e.g. absolute continuity or Hölder regularity). In view of the two-marginal case, one can similarly ask whether there exists a lift $\pi$ that realizes this regularity in the sense that
@@
Similarly, one expects that any alternative lift produces higher energy. This is indeed the case for the regularities previously studied and those in this paper. This observation has motivated introducing 1, where the general functional $\Psi$ is chosen here as $|\cdot|^{p}$. We call a lift minimizing or optimal if it achieves the minimum possible energy. We further call it realizing if it attains equality (1.3) (we will observe that in some cases, a minimizing lift exists but a realizing one doesn't). Lifts that realize the regularity of Wasserstein curves have been previously constructed under different assumptions on regularity, some of which are highlighted here:
@@
    $p$-absolutely continuous curves in $p$-Wasserstein spaces with $p>1$ on separable Hilbert spaces: Ambrosio–Gigli–Savaré [6] characterize these curves via the continuity equation and obtained a vector field $v_{t}$ and a lift $\pi$ on $C(I;\mathcal{X})$ that realizes the metric speed i.e. $|\dot{\mu}_{t}|^{p}=\int_{\mathcal{X}}|v_{t}|^{p}{\mathrm{\,d}}\mu_{t}=\int_{C}|\dot{\gamma}_{t}|^{p}{\mathrm{\,d}}\pi$ for a.e. $t\in I$.
@@
Our study of low-regularity paths on Wasserstein spaces is motivated by measure-valued solutions to conservative stochastic PDEs, such as stochastic Fokker–Planck–Kolmogorov equations. Low-regularity paths can also arise deterministically as solutions to the continuity equation, though less trivially. In this paper, we focus on the deterministic case, and study the random setting in the subsequent companion paper [1]. Since, in low regularity, it is less straightforward which norm can be minimized and potentially realized, our first step is to identify a sufficient condition on the energy to guarantee the existence of a minimizer. Below is Proposition 3.1—a simple observation using the direct method in the calculus of variations.
@@
where $\bar{x}\in\mathcal{X}$ is an arbitrary point and $|\cdot|:C(I;\mathcal{X})\to[0,+\infty]$ is a (semi-)norm. When $(\mathcal{X},d)$ has enough structure so that closed bounded sets are compact, Arzelà-Ascoli theorem can verify which norms $|\gamma|$ make the sublevels of $\Psi$ relatively compact. This is summarized for some commonly used norms in low-regularity settings in Table 1 at the end of this section. We note that the modulus of continuity $w_{\delta}(\gamma)$, $p$-variation $|\gamma|_{p\textrm{-}\mathrm{var}}$, and its infinitesimal characterization $|\gamma|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}$ fail these conditions. In particular, quadratic variation $\gamma]$ in the sense of stochastic analysis also fails; see Remark 1.13. By contrast, Hölder regularity $|\gamma|_{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$, fractional Sobolev regularity $|\gamma|_{W^{\alpha,p}}$, certain Besov regularity $|\gamma|_{b^{\alpha,p}}$, and Sobolev regularity $|\gamma|_{W^{1,p}}$ satisfy the required conditions. Definitions and properties of these norms are given in Sections 2.1, 2.2, 2.3, 2.4 and 2.5. The function space $W^{1,p}(I;\mathcal{X})$ used in this paper coincides with the set of $p$-absolutely continuous curves $AC^{p}(I;\mathcal{X})$; thus, the last item in the table is already addressed by Lisini's results [17].
Following this observation, we consider paths in the fractional Sobolev space $W^{\alpha,p}$ or certain Besov space $b^{\alpha,p}$. These spaces also contain Hölder spaces for specific exponents. Given a metric space $(\mathcal{X},d)$, an integrability parameter $1\leq p<\infty$, and a regularity parameter $0<\alpha<1$, these function spaces are defined as:
@@
Let $(\mathcal{X},d)$ be a complete separable metric space in which closed bounded sets are compact, and $I\coloneqq[0,T]\subset\mathbb{R}$. Let $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ be such that $\mu_{0}\in P_{p}(\mathcal{X})$ and it has a lift with finite $W^{\alpha,p}$-energy with $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, i.e., (1.7) holds for a lift. Then, there exists a minimizer $\pi\in P(C(I;\mathcal{X}))$ to 1 for the energy $\Psi(\gamma)=|\gamma|^{p}_{W^{\alpha,p}}$. In particular,
@@
    Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ can form a compatible collection under suitable conditions on their covariance matrices. See [23, page 49].
@@
    For more nontrivial compatible examples, see [9, Proposition 4.1] and [23, Section 2.3.2].
@@
Figure 1. An illustration of two different constructions, showing which two-dimensional marginals of $\Upsilon_{n}$ are optimal. Here, $n=2$ and the time interval $[0,1]$ is divided into $2^{2}=4$ equal pieces.
@@
$C^{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}\overset{\textrm{trivial}}{\subset}W^{\alpha,p}\,\,\subset C^{\alpha-\frac{1}{p}\textrm{-}\mathrm{H\ddot{o}l}},\qquad W^{\alpha,p}\,\,\subset C^{\frac{1}{\alpha}\textrm{-}\mathrm{var}},$ (1.14)
@@
Another natural question is whether it is possible to remove the assumptions of the existence of a lift with finite energy and compatibility from Theorem 1.3 and Theorem 1.6, respectively, while still obtaining a minimizing lift. We demonstrate that the answer to this question can be negative. In other words, if compatibility is dropped, the infimum (1.1) for the energy $\Psi(\gamma)=|\gamma|^{p}_{W^{\alpha,p}}$ can be $\infty$. More specifically, in Example 4.5, we show:
@@
There exists a non-compatible curve $(\mu_{t})\in W^{\alpha,p}(I;P_{p}(\mathcal{X}))$ for some $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, whose only existing lift $\pi$ on continuous paths has infinite $|\cdot|^{p}_{W^{\alpha,p}}$-energy.
@@
As a final result, we give a dynamic formulation of the Wasserstein distance using Besov energy. We first show in Lemmas 3.11 and 3.12 that on a metric space $(\mathcal{X},d)$ and for any $1<p<\infty$ and $0<\alpha<1$, the following are equivalent:
@@
Notice that $\left(1-2^{-(p-\alpha p)}\right)\in(0,1)$ on the parameter range above. This characterization allows us to give the result below, which is a generalization of the metric Benamou–Brenier formula [7], [6, Eq. (8.0.3)], [17, Corollary 4], to the fractional setting. Below is Corollary 3.13 and $Geo([0,1];\mathcal{X})$ denotes the set of constant-speed geodesics.
@@
Let $(\mathcal{X},d)$ be a complete, separable, and geodesic metric space. Let $1<p<\infty$ and $0<\alpha<1$. Then for every $\mu,\nu\in P_{p}(\mathcal{X})$, we have
@@
In other words, the task is to find a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a stochastic process $(X_{t})$ with marginals $(\mu_{t})$ such that the expectation (1.18) is minimized. As before, if $\mathrm{Process}(\mu_{t})=\varnothing$, the infimum (1.18) is set $\infty$. As in the classical optimal transport, only the distribution of these processes enters the cost function. Thus, we will only focus on the formulation of 1.
@@
Given $\gamma\in C([0,T];\mathcal{X})$, the stochastic calculus quadratic variation $\gamma]_{T}$ over $[0,T]$ is usually defined by taking a limit over shrinking partitions of $[0,T]$. By the example in Remark 2.14, one sees that the map $\gamma\mapsto[\gamma]_{T}$ from $C([0,T];\mathcal{X})\to[0,+\infty]$ is not lower semi-continuous. Furthermore, in Example 4.3, by taking $\upgamma=\tfrac{1}{2}$ and $p>2$, one sees that there exists a determinist curve $(\mu_{t})\subset P_{p}(\mathcal{X})$ with non-zero quadratic variation with respect to $W_{p}$, whose only path-continuous process $(X_{t})$, equivalently its lift $\pi$, has absolutely continuous sample paths. In particular,
$\mu]_{T}^{2}>\mathbb{E}\Big[[X]_{T}^{2}\Big]=\int_{\Gamma_{T}}[\gamma]_{T}^{2}{\mathrm{\,d}}\pi=0,$ (1.20)
which shows that $\mu]_{T}$ cannot even be bounded by the quadratic variation energy of its lift.